# Latent Denoising Makes Good Visual Tokenizers

### v1

@@ 5.3 Benchmarking with Previous Systems
Figure 6: Qualitative Results. We show selected examples of class-conditional generation on ImageNet 256$\times$256 using MAR-L [36] trained with our tokenizer.
@@ 6 Discussion and Conclusion
Limitations. On the general side, our study primarily investigates continuous-valued tokenizers; the effectiveness of our denoising strategy for discrete, vector-quantized (VQ) tokenizers remains an open question. Exploring denoising-based objectives in the context of VQ tokenizers is a valuable future direction. Additionally, our evaluation is currently limited to the ImageNet benchmark. Assessing tokenizer scalability and generalization on broader, real-world datasets is an important next step. On the technical side, we observe a training/inference discrepancy in our tokenizer: the decoder is primarily trained on noise-injected latent embeddings, whereas it operates on almost noise-free embeddings during inference. Fine-tuning the decoder on clean latent embeddings partially addresses this discrepancy. Further investigation into mitigating this discrepancy could yield additional improvements. Nonetheless, the core motivation and insights of our work remain robust.
Conclusion. The strong generalization and effectiveness of our $l$-DeTok across different generative models highlights a fundamental yet underexplored opportunity: tokenizer representations alone can substantially advance different generative approaches without architectural changes. Our findings suggest that explicitly aligning tokenizer training with downstream denoising tasks is surprisingly beneficial for generative modeling, complementing traditional focuses on pixel-level accuracy or semantic alignment. Given the foundational role of tokenizers, we hope this perspective will inspire further research into effective and generalizable representation learning for generative models. Lastly, the core idea of this work is general and extends beyond image tokenization. Exploring its potential in broader generative modeling, such as video generation, action generation, protein design, and other domains, will be an exciting direction for future research.
Acknowledgment
The USC Geometry, Vision, and Learning Lab acknowledges generous supports from Toyota Research Institute, Dolby, Google DeepMind, Capital One, Nvidia, and Qualcomm. Yue Wang is also supported by a Powell Research Award.
@@ A.1 Tokenizer
Our tokenizers are based on ViTs [18]. We provide detailed model parameters in Table A.1. We include the details provided in Section 4 here for completeness: We implement our tokenizer using ViTs for both encoder and decoder [61, 18]. We adopt recent architectural advances from LLaMA [60], including RoPE [57] (together with learned positional embeddings following [22]), RMSNorm [71], and SwiGLU-FFN [53]. The encoder operates at a patch size of $16$, yielding $256$ latent tokens for each $256\times 256$ image, while the decoder uses patch size $1$ as there is no resolution change. The latent dimension is set to $16$. We omit the [CLS] token for simplicity.
@@
For ablation studies, we use a ViT-S encoder and a ViT-B decoder, disable GAN loss, and train for 50 epochs. For our final experiments, we adopt ViT-B for both encoder and decoder, enable GAN loss from epoch 100, and train for 200 epochs. In both settings, the global batch size is 1024, and the peak learning rate is set to $4.0\times 10^{-4}$ (scaled linearly from $1.0\times 10^{-4}$ at a batch size of 256). We apply linear warm-up for 25% of the total epochs (12 epochs for 50-epoch training, and 50 epochs for 200-epoch training), followed by cosine learning rate decay. We use the AdamW optimizer with $\beta$ parameters (0.9, 0.95) and a weight decay of $1.0\times 10^{-4}$. The only data augmentation employed is horizontal flipping. Our reconstruction loss closely follows the implementation in [69].
@@
Decoder fine-tuning.
We observe an interesting discrepancy between training and inference in our $l$-DeTok. Our decoder is trained predominantly with noise-corrupted latent embeddings but encounters nearly clean embeddings during inference. To address this gap, we fine-tune the decoder on clean latent embeddings, i.e., masking and latent noising are disabled, for an additional 100 epochs. This adjustment partially alleviates the discrepancy, improving the FID of the 800-epoch MAR-L model from 1.43 to 1.35 and MAR-B model from 1.61 to 1.55 (Table 2). Figure A.2 and Figure A.1 compare the denoising capability of different tokenizers. The fine-tuned decoder exhibits reduced denoising ability.
Importantly, we find that the performance improvements from our approach primarily result from better latent representations rather than increased denoising capability of the decoder. To validate this, we perform an additional experiment by fine-tuning only the MAR-VAE decoder on the denoising task—keeping the encoder frozen, thus leaving latent representations unchanged. Contrary to expectations, this experiment leads to degraded performance, indicating that merely enhancing the decoder’s denoising capability exacerbates the training-inference discrepancy, since the generative models already excel at this task. These observations confirm that the key advantage of our method lies in improved latent representations, rather than decoder-side denoising improvements.
@@
def denoise(x, encoder, decoder, max_mask_ratio=0.7, gamma=3.0):
# encode input image to latent embeddings under (optional) masking
z, ids_restore = encoder(x, max_mask_ratio=max_mask_ratio)
# variational latent embeddings
posteriors = diagonal_gaussian_dist(z)
z_sampled = posteriors.sample()
# sample interpolation factor uniformly from [0, 1]
bsz, n_tokens, chans = z_sampled.shape
device = z_sampled.device
noise_level = torch.rand(bsz, 1, 1, device=device).expand(-1, n_tokens, chans)
# generate Gaussian noise
noise = gamma * torch.randn(bsz, n_tokens, chans, device=device)
# interpolate latent embeddings with noise
z_noised = (1 - noise_level) * z_quantized + noise_level * noise
# reconstruct the inputs
recon = decoder(z_noised, ids_restore)
return recon
@@ A.2 Generative Models
We re-implement all generative models within our codebase to standardize training and evaluation. We introduce one modification: for DiT [46], SiT [42], and LightningDiT [64], we apply classifier-free guidance (CFG) to all latent channels, rather than only the first three channels used in their original implementations. [^1] See original implementations in DiT, SiT, and LightningDiT. For training on 1D tokens (e.g., MAETok [6]), we use simple 1D learnable positional embeddings and disable RoPE used in LightningDiT. We adopt the default samplers from the original implementations, using 250 denoising steps during sampling.
For autoregressive (AR) methods, we follow the default MAR model and implement RandomAR and RasterAR following RAR [68]. To produce continuous tokens, we employ a diffloss [36] with a 3-layer, 1024-channel MLP head. We disable dropout in all MLP layers within Transformer blocks for autoregressive models. For inference, we use 64 autoregressive steps and 100 denoising steps for all experiments, except those in system-level comparison (Table 2), where we use 256 autoregressive steps. Sampling is performed with the default MAR sampler across all AR models. Following MAR [36], we set the sampling temperature to 1.0 when using CFG, and sweep temperatures when CFG is disabled (i.e., CFG=1.0).
@@
We follow the training recipe from [64] for all experiments reported in this paper: global batch size of 1024, AdamW optimizer, constant learning rate of $2\times 10^{-4}$, and no warm-up, gradient clipping, or weight decay. We do not tune these hyperparameters.

### v2

@@ 5.5 Beyond Standard Continuous Tokenizers
Figure 6: Qualitative Results. We show selected examples of class-conditional generation on ImageNet $\times256$ using MAR-L (Li et al., 2024a) trained with our tokenizer.
@@ 6 Discussion and Conclusion
Simple principles that scale well are at the core of generative modeling. In this work, we have shown that a surprisingly simple principle, i.e., denoising-based tokenizer, already provides consistent benefits across a broad design space, including non-autoregressive and autoregressive generative models, 2D/1D/VQ tokenizers, CNN and Transformer-based architectures, and both raster and random generation orders, while adding almost no system complexity. Several questions remain open.
First, our experiments are conducted within a moderate-scale regime ($\mathcal{O}(10^{6})$ samples), where $l$-DeTok improves performance when we scale up within this envelope. A natural next step is to test it in the frontier regime of web-scale models and datasets ($\mathcal{O}(10^{7})$$\mathcal{O}(10^{10})$ samples), more challenging settings such as video generation, and to understand how denoising-aligned tokenizers behave in those cases. Second, our results also parallel the recent trend of semantics distillation from pre-trained encoders such as DINOv2. Distillation is effective when strong teachers exist and align well with the target domain, but it may become a bottleneck as we move to larger data and broader distributions that surpass the teacher’s coverage, or to domains without reliable teachers. In contrast, $l$-DeTok is self-contained and does not depend on external teachers. At scale, we hope such task-aligned tokenizers may provide a more flexible path forward. Finally, reconstructing clean inputs from noisy latents makes our tokenizer objective reminiscent of $x_{0}$-prediction in modern generative models. Understanding the boundary—and eventual unification—among reconstruction, denoising and generation is an intriguing direction, e.g., Li and He (2025). We hope this perspective will help guide future work on scalable generative models and the tokenizers that support them.
@@ A.1 Tokenizer
We implement our tokenizer using ViTs for both encoder and decoder (Vaswani et al., 2017; Dosovitskiy et al., 2021). We provide detailed model parameters in Table A.1. We adopt recent architectural advances from LLaMA (Touvron et al., 2023), including RoPE (Su et al., 2024) (together with learnable positional embeddings following Fang et al. (2023)), RMSNorm (Zhang and Sennrich, 2019), and SwiGLU-FFN (Shazeer, 2020). The encoder operates at a patch size of $16$, yielding $256$ latent tokens for each $256\times 256$ image, while the decoder uses patch size $1$ as there is no resolution change. The latent dimension is set to $16$. We omit the [CLS] token for simplicity.
@@
Our tokenizer is trained using the weighted loss defined in Eq. 6, with default weights $\lambda_{\text{KL}}=10^{-6}$, $\lambda_{\text{percep}}=1.0$, and $\lambda_{\text{GAN}}=0.1$. For ablation studies, we use a ViT-S encoder and a ViT-B decoder, disable GAN loss, and train for 50 epochs. We observe that including a GAN loss (Yu et al., 2021; Goodfellow et al., 2014; Esser et al., 2021) sharpens reconstructions but roughly doubles training time, without altering result trends. For our final experiments, we adopt ViT-B for both encoder and decoder, enable GAN loss from epoch 100, and train for 200 epochs. In both settings, we use a global batch size of 1024, and the peak learning rate is set to $4.0\times 10^{-4}$ (scaled linearly from $1.0\times 10^{-4}$ at a batch size of 256 (Goyal et al., 2017)). We apply linear warm-up for 25% of the total epochs (12 epochs for 50-epoch training, and 50 epochs for 200-epoch training), followed by cosine learning rate decay. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with $\beta$ parameters (0.9, 0.95) and a weight decay of $1.0\times 10^{-4}$. The only data augmentation employed is horizontal flipping. Our reconstruction loss closely follows the implementation in Yu et al. (2025a).
Training budget.
@@
We present a detailed comparison of training cost between our baseline and $l$-DeTok tokenizers in Table A.2. All configurations use the same architecture, batch size (1024), and total number of epochs; the only difference is whether we use the baseline objective or $l$-DeTok. As shown, the approximate A100 hours of our method are essentially identical to, or slightly lower than, the baseline: S-B goes from 180 (baseline) to 160–180 hours with $l$-DeTok, B-B from 1150 to 1140 hours, and the 1D and VQ variants remain within the same 170–190 hour range. This indicates that the denoising mechanism does not introduce any meaningful extra computational overhead beyond the standard tokenizer training; in practice, the variation is within normal run-to-run fluctuation.
Table A.2: Training Cost. Comparison of training schedules and approximate compute for different tokenizer and setting combinations. Our models are trained across mixed compute and storage environments (e.g., A100s, H100s, and A6000s). We report approximate A100-equivalent hours over the full training run. Size: Encoder size - Decoder size. S: Small, B: Base.
Size Setting Epochs Batch size GAN start epoch Approx. A100 Hours
S-B Baseline 50 1024 disabled 180
S-B $l$-DeTok 50 1024 disabled 160–180
B-B Baseline 200 1024 100 1150
B-B $l$-DeTok 200 1024 100 1140
B-B $l$-DeTok + Distill. 200 1024 100 1900
S-B 1D-baseline 50 1024 disabled 170–190
S-B 1D-$l$-DeTok 50 1024 disabled 170–190
S-B VQ-baseline 50 1024 disabled 190
S-B VQ-$l$-DeTok 50 1024 disabled 170
@@ B-B $l$-DeTok 200 1024 100 1140
1def denoise(x, encoder, decoder, max_mask_ratio=0.7, gamma=3.0):
2 # encode input image to latent embeddings under (optional) masking
3 z, ids_restore = encoder(x, max_mask_ratio=max_mask_ratio)
4
5 # variational latent embeddings
6 posteriors = diagonal_gaussian_dist(z)
7 z_sampled = posteriors.sample()
8
9 # sample interpolation factor uniformly from [0, 1]
10 bsz, n_tokens, chans = z_sampled.shape
11 device = z_sampled.device
12 noise_level = torch.rand(bsz, 1, 1, device=device).expand(-1, n_tokens, chans)
13
14 # generate Gaussian noise
15 noise = gamma * torch.randn(bsz, n_tokens, chans, device=device)
16
17 # interpolate latent embeddings with noise
18 z_noised = (1 - noise_level) * z_quantized + noise_level * noise
19
20 # reconstruct the inputs
21 recon = decoder(z_noised, ids_restore)
22 return recon
@@ A.2 Generative Models
To evaluate the broad effectiveness of a tokenizer, we experiment with six representative generative models, including three non-autoregressive models: DiT (Peebles and Xie, 2023), SiT (Ma et al., 2024), and LightningDiT (Yao and Wang, 2025); and three autoregressive models: MAR (Li et al., 2024a), causal RandomAR, and RasterAR based on RAR (Yu et al., 2024) and diffloss (Li et al., 2024a). We follow their officially released code to re-implement all generative models within our codebase to standardize training and evaluation, ensuring fair comparisons.
We introduce some modifications: for DiT (Peebles and Xie, 2023), SiT (Ma et al., 2024), and LightningDiT (Yao and Wang, 2025), we apply classifier-free guidance (CFG) to all latent channels, rather than only the first three channels used in their original implementations. [^2] See original implementations in DiT, SiT, and LightningDiT. For training on 1D tokens (e.g., MAETok (Chen et al., 2025a)), we use simple 1D learnable positional embeddings and disable RoPE used in LightningDiT. We adopt the default samplers from the original implementations, using 250 denoising steps during inference. Due to the frequently observed training instability in SiT and LightningDiT in our early experiments, we apply QK-Norm (Dehghani et al., 2023) to stabilize training. These modifications are consistent across all tokenizers for fairness.
For autoregressive (AR) methods, we use the default MAR model and implement RandomAR and RasterAR following RAR (Yu et al., 2024). To produce continuous tokens, we employ the diffloss (Li et al., 2024a) with a 3-layer, 1024-channel MLP head. We disable dropout in all MLP layers within Transformer blocks for autoregressive models. For inference, we use 64 autoregressive steps and 100 denoising steps for all experiments, except those in system-level comparison (Tables 4 and 5), where we use 256 autoregressive steps or 512 autoregressive steps. Sampling is performed with the default MAR sampler across all AR models. Following MAR (Li et al., 2024a), we set the sampling temperature to 1.0 when using CFG, and sweep temperatures when CFG is disabled (i.e., CFG=1.0).
@@
We use a standardized training recipe for all generative models. Specifically, we follow the hyperparameters from Yao and Wang (2025), training generative models with a global batch size of 1024, using AdamW optimizer (Loshchilov and Hutter, 2019) with a constant learning rate of $2\times 10^{-4}$, without warm-up, gradient clipping, or weight decay. We do not tune these hyperparameters. For ablation studies, we train generative models for 100 epochs. For larger-scale experiments on MAR models, we train them for 800 epochs. All models utilize exponential moving average (EMA) with a decay rate of 0.9999.
Standardization.
We always standardize tokenizer outputs by subtracting the channel-wise mean and dividing by the channel-wise standard deviation, both computed from the ImageNet training set. For publicly available tokenizers, we use their official standardization steps.
Metrics.
Unless noted otherwise, we report FID@50k scores with classifier-free guidance (CFG), with optimal CFG scales searched from FID@10k results. We abbreviate FID@50k as FID.
Training budget.
@@
A.3 Tokenizer tuning
Decoder fine-tuning.
We observe an interesting discrepancy between training and inference in our $l$-DeTok. Our decoder is trained predominantly with noise-corrupted latent embeddings but encounters nearly clean embeddings during inference. To address this gap, we fine-tune the decoder on clean latent embeddings, i.e., masking and latent noising are disabled, for an additional 100 epochs. This adjustment partially alleviates the discrepancy, improving the FID of the 800-epoch MAR-L model from 1.43 to 1.35 and MAR-B model from 1.61 to 1.55 (Table 4). Figure A.1 and Figure A.2 compare the denoising capability of different tokenizers. Although fine-tuning decoder reduces the decoder’s denoising strength, it enhances the image quality (FID and IS) when decoding from generated latent embeddings. Crucially, this demonstrates that the quality of latent representations produced by the tokenizer encoder, rather than the decoder’s error-tolerance, is the primary driver of generative model improvements. Therefore, our $l$-DeTok does more than improving the the decoder’s robustness to sampling errors.
To further verify that performance gains primarily stem from improved latent representations rather than enhanced decoder-side error tolerance, we conduct an additional experiment by fine-tuning only the MAR-VAE decoder on the denoising task while keeping the encoder frozen. In this way, we can reuse pre-trained models for comparison since their modeling space, encoder latent space, remains fixed. Contrary to expectations, this setup—improving decoder robustness without altering latent embeddings—actually leads to slightly worse FIDs (around 0.1 to 0.3 degradation). This indicates that merely boosting decoder denoising capability exacerbates the training-inference discrepancy, as generative models are already proficient at denoising. These results reinforce that the primary advantage of our approach lies explicitly in the improved quality of latent embeddings, not in decoder-side denoising enhancements.
Adopting tokenizers to $512\times 512$ resolution.
To save computation, we fine-tune tokenizer checkpoints from ImageNet-$256\times 256$ for the ImageNet-$512\times 512$ generation task. Specifically, we halve the learning rate to $2.0\times 10^{-4}$ and fine-tune tokenizers initialized from the 200-epoch $256\times 256$ checkpoints for an additional 25 epochs, enabling GAN loss from the start. We compare full fine-tuning to decoder-only fine-tuning: decoder-only fine-tuning achieves better PSNR (26.3 vs. 24.9) and comparable rFID (0.92 vs. 0.86), yet fully fine-tuned tokenizers consistently provide superior downstream generative performance. For instance, using fully fine-tuned tokenizers, MAR-B achieves 2.66 FID and SiT achieves 5.98 FID, compared to decoder-only fine-tuned tokenizers at 3.63 and 6.32 FID, respectively. All generative models are trained from scratch for 100 epochs, and we report FID@50K at optimal CFG scales. Consequently, we adopt fully fine-tuned tokenizers.
These results reinforce our core motivation: encoder-produced tokenizer embeddings primarily drive generative quality, while decoder fine-tuning provides incremental improvements. Therefore, improving encoder is more fundamental than improving the decoder.
Adapting generative models to $512\times 512$ resolution.
Similar to tokenizer fine-tuning, we fine-tune a MAR-L checkpoint from ImageNet-$256\times 256$ for the ImageNet-$512\times 512$ generation task to reduce compute. Additionally, we train a MAR-B model from scratch to verify the compatibility of our tokenizer for end-to-end training. Note that we do not aim to hill climb the performance metrics in Table 5, since our tokenizer is only fine-tuned for 25 epochs rather than trained from scratch for many epochs. Instead, our intention is for these results to demonstrate the broad applicability of our approach.
A.4 $l$-DeTok with Semantics Distillation
Our $l$-DeTok already holds great promise for generative models across both autoregressive (AR) and non-autoregressive (non-AR) methods without semantics distillation. It is natural to question whether incorporating semantics distillation could further improve its performance.
To investigate this, we introduce an auxiliary semantics-distillation loss into our tokenizer training pipeline. Specifically, we implement the latent representation projector as a three-layer MLP, similar to REPA (Yu et al., 2025b). The noised latent embeddings (z_noised at line 18 in Algorithm 1) are projected through the following layers: Linear(16, 2048) $\rightarrow$ SiLU() $\rightarrow$ Linear(2048, 2048) $\rightarrow$ SiLU() $\rightarrow$ Linear(2048, 768), where 16 represents the latent embedding dimension and 2048 the intermediate projection dimension. This projector is trained to maximize the cosine similarity between the projected embeddings and semantic features extracted from the pretrained DINOv2-Base model (Oquab et al., 2024).
We have also experimented with a decoder-based semantics-distillation approach, i.e., using an auxiliary Transformer-based decoder to predict pretrained features. After training both implementations for 50 epochs, we found their results were very similar, though the decoder-based implementation required more computation because of the decoder overhead. Thus, we adopted the simpler MLP-based approach for efficiency.
The loss coefficient for minimizing the cosine distance is set to $1.0$ for simplicity, which we do not sweep or tune. We train this tokenizer using the identical settings as the “+Extended” protocol described in Table 5: 200 epochs in total, activating GAN loss from epoch 100.
The results from this experiment (summarized in Table 3) demonstrate substantial improvements, particularly for diffusion-based non-AR models (e.g., SiT). With semantics distillation, our $l$-DeTok achieves state-of-the-art FID scores for non-AR models, clearly surpassing previous semantics-distilled tokenizers such as MAETok (Chen et al., 2025a) and VA-VAE (Yao and Wang, 2025). While AR models benefit relatively less from distillation—and performance may even slightly degrade compared to the non-distilled $l$-DeTok—their generative quality remains notably better than previous approaches.
Nonetheless, we emphasize the practical significance of our original $l$-DeTok formulation without semantics distillation, especially for domains like video, audio, proteins, poses, or trajectories, where strong pretrained semantic models comparable to DINOv2 (Oquab et al., 2024) may be unavailable or significantly weaker. Thus, our proposed tokenizer remains broadly valuable both with and without semantic distillation.
A.5 Convolutional Tokenizer
Model.
We reuse MAR-VAE and SD-VAE architectures with a downsampling ratio of 16 and latent dimension 16, producing latent representations of shape $(16,16,16)$ for $(C,H,W)$. We add only a few lines of latent-denoising code (similar to Algorithm 1) to CNN-based tokenizers. Tokenizers are trained for 50 epochs, both with and without latent-denoising, using hyperparameters identical to our Transformer-based tokenizers, potentially disadvantaging CNN models. [^3] In our experience, Transformer- and CNN-based tokenizers typically require distinct hyperparameter sets. We enable GAN loss from epoch 10 for CNN tokenizers. We do not sweep the noise standard deviation and only run with $\gamma=3.0$ and $\gamma=0.0$ (baseline). Further hyperparameters and noise strength tuning and longer training may yield additional improvements for CNN-based tokenizers, which we leave for future work.
Results.
We report FID@50k under best CFG scales. With the default CNN baseline, MAR-B achieves 3.32 FID and SiT-B 7.11 FID. With our denoising-based CNN tokenizers, MAR-B improves to 2.82 FID and SiT-B to 5.62 FID, demonstrating that $l$-DeTok provides consistent gains also in convolutional networks. Our results clearly demonstrate the general, architecture-agnostic nature of our latent-denoising idea.
(a) Baseline 1D tokenizers.
(b) Naive denoising with $\gamma=2.0$ (delayed “grokking”).
(c) Denoising with $\gamma=2.0$ enabled after warm-up.
Figure A.3: Training dynamics of 1D tokenizers. PSNR curves for different token counts $K$. (a) Baseline 1D tokenizers exhibit a “grokking”-like phase where PSNR suddenly improves after a plateau. (b) Directly adding denoising with $\gamma=2.0$ delays this transition, especially for larger $K$. (c) Enabling denoising only after the baseline grokking phase (10-epoch warm-up) yields stable training across all $K$. X-axis: calibrated global step, 1k-step denotes 1 ImageNet epoch.
A.6 1D Tokenizers
Implementations. We construct 1D continuous tokenizers by minimally modifying the 2D continuous tokenizers used in our main experiments. Concretely, we keep the image encoder and decoder backbones unchanged and only alter the latent parameterization and positional embeddings. Following Yu et al. (2025a), we append $K\in\{32,64,128\}$ learnable latent tokens to the encoder input sequence, leading to $256+K$ tokens for a $256\times 256$ image with patch size 16 (i.e., 256 patch tokens plus $K$ latent tokens). After encoding, only the $K$ latent tokens are passed to the decoder. The decoder then uses learnable mask tokens to reconstruct the original images from these $K$ latents. All Transformer blocks are based on self-attention. For simplicity, we drop RoPE and use only learnable absolute positional embeddings.
1D tokenizers with $l$-DeTok. When applying $l$-DeTok to 1D tokenizers, we use interpolative latent noising with $\gamma=2.0$ and omit masking entirely for simplicity. All other training hyperparameters (optimizer, schedule, batch size, and loss) are kept identical to those used for our 2D continuous tokenizers, which is convenient for comparison but may be suboptimal for the 1D tokenizers.
Grokking. We observe a practical “grokking”-like phase when training 1D tokenizers: AA (Fig. A.3a). This behavior already appears in the baseline 1D tokenizers without $l$-DeTok. A naive application of denoising with $\gamma=2.0$ (without any warm-up) further delays this transition, especially for larger $K$ (Fig. A.3b), so under a fixed training budget the tokenizers remain undertrained and yield poor downstream FID. To mitigate this, we warm up the 1D tokenizer for 10 epochs by disabling the denoising task, and then enable it from epoch 10 onward with a milder noising level $\gamma=2.0$ to reduce task difficulty. [^4] We initially used $\gamma=3.0$ and observed poor downstream performance. We first attributed this failure to the large $\gamma$, but later found that the delayed grokking was the primary cause. This staged schedule makes all runs stable and successful (Fig. A.3c), and is analogous to common tokenizer training practices that warm up the model before turning on GAN or perceptual losses. We hope this full transparency will serve as a useful reference and provide concrete data points for future research.
A.7 Vector-quantized (VQ) Tokenizers
Implementations. For the VQ tokenizer, we start from the same continuous tokenizer architecture and only replace the bottleneck head. In the continuous variant, a linear layer maps Transformer tokens from the model width to the latent channels (e.g., $16$ or $2\times 16$ for the variational autoencoder). In the VQ variant, this head is replaced by a VQ bottleneck: a linear projection first maps each token to a 64-dimensional latent vector, which is then quantized using a codebook of size 4096, i.e., a learnable matrix of shape $(4096,64)$. Following prior work (Yu et al., 2021; 2025a), we $\ell_{2}$-normalize the codebook vectors and train with the standard VQ objective (reconstruction loss plus codebook loss) and a commitment loss with weight 0.25; we refer readers to these works for further formulation details. All other components of the encoder and decoder are unchanged. As with our 1D and continuous tokenizers, we reuse the same training hyperparameters (optimizer, schedule, and loss weights apart from the VQ terms), which simplifies comparison but may be suboptimal for the VQ setting.
Generators. On top of the learned VQ tokenizer, we train discrete auto-regressive (AR) generators in two settings: (i) RandomAR, where the 2D grid of discrete tokens is visited in a random permutation at both training and sampling time; and (ii) RasterAR, where the same architecture is trained with the standard raster-scan ordering. We directly use the implementation from RAR (Yu et al., 2024) without modification and fix the generation order (random or raster-scan) consistently during both training and inference. These models are decoder-only, causal-attention Transformers, trained with cross entropy loss in a teacher-forcing manner.
VQ tokenizers with $l$-DeTok. Our $l$-DeTok variant for VQ tokenizers is intentionally minimal: we only leverage masking as deconstruction during training. Starting from the baseline RasterAR setup, we sweep masking ratios of 0.7 and 0.95. Both ratios yield better downstream FID than the baseline VQ tokenizer, with 0.95 performing best. We therefore adopt a masking ratio of 0.95 for RandomAR runs.